# Quantum speed limit in quantum sensing

### v1

@@ Abstract
Quantum sensors capitalize on advanced control sequences for maximizing sensitivity and precision. However, protocols are not usually optimized for temporal resolution. Here, we establish the limits for time-resolved sensing of dynamical signals using qubit probes. We show that the best possible time resolution is closely related to the quantum speed limit (QSL), which describes the minimum time needed to transform between basis states. We further show that a bipartite control sequence consisting of two phase-shifted pulses reaches the QSL. Practical implementation is discussed based on the example of the spin-1 qutrit of a nitrogen-vacancy (NV) center in diamond.
The energy-time uncertainty principle introduced by Heisenberg is a fundamental concept of quantum mechanics. While formulated loosely in Heisenberg’s initial work, Robertson [1] and Bohr [2] put it on firm ground by formalizing the relationship between uncertainty and non-commutativity of observables. Twenty years later, Mandelstam and Tamm [3] showed that the energy-time uncertainty is not an uncertainty relation due to non-commutativity, but rather a statement about intrinsic time scales of quantum systems [4]. This insight led to the derivation of a quantum speed limit (QSL). To do so, Mandelstam and Tamm used the von Neumann equation with the projection operator to develop an expression for the overlap between time-evolved states,
@@
where $\langle\Delta H\rangle^{2}=\langle H^{2}\rangle-\langle H\rangle^{2}$ is the variance of the Hamiltonian. The minimum time needed to obtain a fully orthogonal state, i.e., $\braket{\psi(0)}{\psi(t)}=0$, is then given by
@@
Since the initial formulations of the QSL, theoretical work has focused on more complex situations such as open quantum systems [9] and systems with quantum entanglement [10], as well as applications in quantum information processing [11]. Here, the QSL sets the maximum speed at which computations can be performed and therefore permits deriving an upper bound of the computational limits of the universe [12, 13]. Further, the QSL has been studied in the field of quantum optimal control [14], quantum thermodynamics [15], and metrology with respect to quantum clocks [16]. Obviously, the finite time of qubit operations will also limit the time resolution achievable in quantum sensing tasks [17]. The time limitation is of practical importance because of the many existing and envisioned applications of quantum sensors.
In this work, we investigate the relation between the QSL and the time resolution achievable in quantum sensing experiments. We show that a bipartite sensing sequence consisting of two phase-shifted control rotations, equivalent to a Ramsey sequence with zero time delay, reaches the QSL. Opposite to a Ramsey measurement, however, phase accumulation occurs during control rotations rather than a free evolution interval. We derive quantitative expressions for the quantum phase pick-up as a function of control rotation angle and velocity. We use these expressions to define the time resolution and bandwidth of the sensing sequence. We also show that time resolution can be extended beyond the QSL by trading for a reduced signal-to-noise ratio. As an example, we simulate the expected response for a nitrogen-vacancy (NV) center in diamond that is exposed to rapidly varying magnetic signal.
@@
Ideally, the control rotations $\hat{R}$ and $\hat{R}^{\prime}$ are infinitely fast [Fig. 1(c)]. However, owing to the QSL, $\hat{R}$ and $\hat{R}^{\prime}$ must have a finite duration $t_{\mathrm{R}}$ or, equivalently, a finite angular velocity $\Omega$. In experiments, the maximum $\Omega$ may be limited by many factors, including finite available driving power, competition between $\omega_{0}$ and $\Omega$, excitation of further lying energy levels, or a combination of those. The finite speed of control rotations fundamentally limits the time resolution of the sensing sequence.
For finite $t_{\mathrm{R}}$, qubit evolution under the signal Hamiltonian already occurs during the control rotations. The qubit-acquired phase then becomes the sum of the phase pick-up during the free evolution interval and the phase pick-up during control rotations [Fig. 1(d)]. In the most extreme case, the free evolution time is zero, and phase pick-up occurs entirely during control rotations [Fig. 1(e)]. In this situation, the time resolution of the sequence reaches that of the QSL.
We next derive quantitative relationships between the phase pick-up $\phi_{\mathrm{eff}}$ and the velocity $\Omega$ and angle $\alpha$ of control rotations. For our derivation, we focus on a bipartite control sequence consisting of two consecutive qubit rotations $\hat{R}=\hat{R}_{Y}^{\alpha}$ and $\hat{R}^{\prime}=\hat{R}_{\pm X}^{\alpha}$ of equal duration $t_{\mathrm{R}}$ [Fig. 1(e)]. Here, $X$ and $Y$ are orthogonal axes, as shown in Fig. 1(f). This sequence represents the basic Ramsey scheme with zero time delay. The sequence duration is $\tau=2t_{\mathrm{R}}=\alpha/\Omega$, where $\Omega$ is the maximum allowed rotation velocity. In the Supplemental Material [18], we show that an equal time-share ($t_{\mathrm{R}}=t_{\mathrm{R^{\prime}}}$) between the two parts of the bipartite pulse sequence yields optimum sensitivity (as defined by Eq. (8) below). This is also true when varying the angle between rotation axes.
@@
where FT is the Fourier transform.
@@
As a practical example, we consider the $S=1$ spin system of a single nitrogen-vacancy (NV) center in diamond [21]. The NV center is a prototypical qubit sensor with a growing range of applications in materials science, physics, chemistry, and biology [22, 23, 24, 25], including the study of dynamical excitations in these systems [26, 27, 28]. The NV center exhibits three spin energy levels $m_{S}=0$ and $m_{S}=\pm 1$ and two allowed spin transitions with frequencies $\omega$ and $\omega^{\prime}$, as depicted in Fig. 4(a). To form an effective two-level system, the $\omega$ (or $\omega^{\prime}$) transition can be isolated by applying an axial bias field $B_{0}$ along the NV symmetry axis [Fig. 4(b)]. For driving fields $\Omega$ smaller than the frequency difference $|\omega^{\prime}-\omega|$ between the $m_{S}=\pm 1$ levels [Fig. 4(c)], the excitation is selective and the other ($\omega^{\prime}$ or $\omega$) transition can be neglected.
@@
In summary, our work establishes limits to the temporal resolution reachable by a quantum sensor. We develop our discussion in the framework of the QSL, which is applied to the canonical sensing principle of Ramsey interferometry. We derive expressions for the coherent phase pick-up during a generic sequence consisting of two phase-shifted control pulses and analyze the response regarding time resolution, frequency bandwidth, and sensitivity. We numerically simulate the expected phase response for the single spin of an NV center in diamond, taking the full $S=1$ nature of the spin system and nonlinear driving effects into account. Beyond fundamental aspects in quantum metrology, our work has practical relevance for real-world applications of quantum sensors, such as the mapping of fast magnetization or charge dynamics in materials and integrated devices.

### v2

@@ Abstract
Quantum sensors capitalize on advanced control sequences for maximizing sensitivity and precision. However, protocols are not usually optimized for temporal resolution. Here, we establish the limits for time-resolved sensing of dynamical signals using qubit probes. We show that the best possible time resolution is closely related to the quantum speed limit (QSL), which describes the minimum time needed to transform between basis states. We further show that a composite control sequence consisting of two phase-shifted pulses reaches the QSL. Practical implementation is discussed based on the example of the spin-1 qutrit of a nitrogen-vacancy (NV) center in diamond.
The energy-time uncertainty principle introduced by Heisenberg is a fundamental concept of quantum mechanics. While formulated loosely in Heisenberg’s initial work, Robertson [1] and Bohr [2] put it on firm ground by formalizing the relationship between uncertainty and non-commutativity of observables. Twenty years later, Mandelstam and Tamm [3] showed that the energy-time uncertainty is not an uncertainty relation due to non-commutativity, but rather a statement about intrinsic time scales of quantum systems [4]. This insight led to the derivation of a quantum speed limit (QSL). To do so, Mandelstam and Tamm used the von Neumann equation with the projection operator to develop an expression for the overlap between the initial state |$\ket{\psi(0)}$ and the time-evolved state |$\ket{\psi(t)}$, yielding
@@
in the domain $0\leq t\leq t_{\mathrm{QSL}}^{(\mathrm{MT})}$ where $\langle\Delta H\rangle^{2}=\langle H^{2}\rangle-\langle H\rangle^{2}$ is the variance of the Hamiltonian. The minimum time needed to obtain a fully orthogonal state, i.e., $\braket{\psi(0)}{\psi(t)}=0$, is then given by
@@
Since the initial formulations of the QSL, theoretical work has focused on more complex situations such as open quantum systems [9] and systems with quantum entanglement [10], as well as applications in quantum information processing [11]. Here, the QSL sets the maximum speed at which computations can be performed and therefore permits deriving an upper bound of the computational limits of the universe [12, 13]. Further, the QSL has been studied in the field of quantum optimal control [14, 15], quantum thermodynamics [16], and metrology with respect to quantum clocks [17]. Obviously, the finite time of qubit operations will also limit the time resolution achievable in quantum sensing tasks [18]. The time limitation is of practical importance because of the many existing and envisioned applications of quantum sensors.
In this work, we investigate the relation between the QSL and the time resolution achievable in quantum sensing experiments. We show that a composite pulse sequence consisting of two phase-shifted control rotations, equivalent to a Ramsey sequence with zero time delay, reaches the QSL. Opposite to a Ramsey interferometry measurement, however, phase accumulation occurs during control rotations rather than a free evolution interval. We derive quantitative expressions for the quantum phase pick-up as a function of control rotation angle and velocity. We use these expressions to define the time resolution and bandwidth of the sensing sequence. We also show that time resolution can be extended beyond the QSL by trading for a reduced signal-to-noise ratio. As an example, we simulate the expected response for a nitrogen-vacancy (NV) center in diamond that is exposed to rapidly varying magnetic signal.
@@
Ideally, the control rotations $\hat{R}$ and $\hat{R}^{\prime}$ are infinitely fast [Fig. 1(c)]. However, owing to the QSL, $\hat{R}$ and $\hat{R}^{\prime}$ must have a finite duration $t_{\mathrm{R}}$ or, equivalently, a finite angular velocity (Rabi frequency $\Omega$). In experiments, the maximum $\Omega$ may be limited by many factors, including finite available driving power, competition between $\omega_{0}$ and $\Omega$, excitation of further lying energy levels, or a combination of those. The finite speed of control rotations fundamentally limits the time resolution of the sensing sequence.
For finite $t_{\mathrm{R}}$, qubit evolution under the signal Hamiltonian already occurs during the control rotations. The qubit-acquired phase then becomes the sum of the phase pick-up during the free evolution interval and the phase pick-up during control rotations [Fig. 1(d)]. In the most extreme case, the free evolution time is zero, and phase pick-up occurs entirely during control rotations [Fig. 1(e)]. In this situation, the time resolution of the sequence reaches that of the QSL defined by Eqs. (2,3). (Note that for a qubit system, the two definitions are equivalent.)
We next derive quantitative relationships between the phase pick-up $\phi_{\mathrm{eff}}$, the velocity $\Omega$ and the angle $\alpha$ of control rotations. For our derivation, we focus on a bipartite control sequence consisting of two consecutive qubit rotations $\hat{R}=\hat{R}_{Y}^{\alpha}$ and $\hat{R}^{\prime}=\hat{R}_{\pm X}^{\alpha}$ of equal duration $t_{\mathrm{R}}$ [Fig. 1(e)]. Here, $X$ and $Y$ are orthogonal axes, as shown in Fig. 1(f). This sequence represents the basic Ramsey scheme with zero time delay. The sequence duration is $\tau=2t_{\mathrm{R}}=\alpha/\Omega$, where $\Omega$ is the maximum allowed rotation velocity. In the Supplemental Material [19], we show that an equal time-share ($t_{\mathrm{R}}=t_{\mathrm{R^{\prime}}}$) between the two parts of the bipartite pulse sequence yields optimum sensitivity (as defined by Eq. (8) below). This is also true when varying the angle between rotation axes.
@@
where FT is the Fourier transform. $K(\omega)$ for other sensing sequences, such as those using amplitude-shaped pulses, can be computed numerically using spin dynamics simulations [19]. Amplitude shaping will, however, always worsen time resolution, because qubit rotations are below the maximum possible $\Omega$.
@@
As a practical example, we consider the $S=1$ spin system of a single nitrogen-vacancy (NV) center in diamond [22]. The NV center is a prototypical qubit sensor with a growing range of applications in materials science, physics, chemistry, and biology [23, 24, 25, 26, 27, 28], including the study of dynamical excitations in these systems [29, 30, 31]. The NV center exhibits three spin energy levels $m_{S}=0$ and $m_{S}=\pm 1$ and two allowed spin transitions with frequencies $\omega$ and $\omega^{\prime}$, as depicted in Fig. 4(a). To form an effective two-level system, the $\omega$ (or $\omega^{\prime}$) transition can be isolated by applying an axial bias field $B_{0}$ along the NV symmetry axis [Fig. 4(b)]. For driving fields $\Omega$ smaller than the frequency difference $|\omega^{\prime}-\omega|$ between the $m_{S}=\pm 1$ levels [Fig. 4(c)], the excitation is selective and the other ($\omega^{\prime}$ or $\omega$) transition can be neglected.
@@
In summary, our work establishes limits to the temporal resolution reachable by a quantum sensor. We develop our discussion in the framework of the QSL, which is applied to the canonical sensing principle of Ramsey interferometry. We derive expressions for the coherent phase pick-up during a generic sequence consisting of two phase-shifted control pulses and analyze the response regarding time resolution, frequency bandwidth, and sensitivity. We numerically simulate the expected phase response for the single spin of an NV center in diamond, taking the full $S=1$ nature of the spin system and nonlinear driving effects into account. Beyond fundamental aspects in quantum metrology, our work has practical relevance for real-world applications of quantum sensors, such as the mapping of fast magnetization reversals in spintronic devices. These dynamics typically occur on time scales of a few nanoseconds  [35]. For NV centers, $t_{\mathrm{min}}\sim 1\,\mathrm{ns}$ is reached, for example, using $\Omega/2\pi=100\,\mathrm{MHz}$ and $\alpha=60^{\circ}$ [Eq. (13)]. Such Rabi frequencies $\Omega$ are available using on-chip microwave delivery [33]. An initial experimental demonstration of our technique is discussed in Ref. [36].